# Mechanistic Interpretability for AI Safety A Review

### v1

@@ Abstract
Understanding AI systems’ inner workings is critical for ensuring value alignment and safety. This review explores mechanistic interpretability: reverse-engineering the computational mechanisms and representations learned by neural networks into human-understandable algorithms and concepts to provide a granular, causal understanding. We establish foundational concepts such as features encoding knowledge within neural activations and hypotheses about their representation and computation. We survey methodologies for causally dissecting model behaviors and assess the relevance of mechanistic interpretability to AI safety. We investigate challenges surrounding scalability, automation, and comprehensive interpretation. We advocate for clarifying concepts, setting standards, and scaling techniques to handle complex models and behaviors and expand to domains such as vision and reinforcement learning. Mechanistic interpretability could help prevent catastrophic outcomes as AI systems become more powerful and inscrutable.
@@ 1 Introduction
The structure of this paper provides a cohesive overview of mechanistic interpretability, situating the mechanistic approach in the broader interpretability landscape (Section 2), presenting core concepts and hypotheses (Section 3), explaining methods and techniques (Section 4), discussing evaluation (Section 5), presenting a taxonomy and survey of the current field (Section 6), exploring relevance to AI safety (Section 7), and addressing challenges (Section 8) and future directions (Section 9).
@@ 2 Interpretability Paradigms from the Outside In
interpretability adopts a top-down approach to unraveling a model’s decision-making processes by probing its learned representations for high-level concepts and patterns governing behavior. Techniques in this paradigm range from training supervised auxiliary classifiers (Belinkov, 2021) to employing unsupervised contrastive and structured probes (see Section 4.1) that explore the model’s latent knowledge (Burns et al., 2023). Beyond observational analysis, concept-based interpretability can enable manipulation of these representations (representation engineering (Zou et al., 2023)), potentially enhancing safety by upregulating concepts such as honesty, harmlessness, and morality.
@@ 3 Core Concepts and Assumptions
This section introduces the foundational concepts and hypotheses that underpin mechanistic interpretability, including the notion of features as fundamental units of representation and their computation through circuits (Section 3.1), and the implications of these concepts for understanding the emergent properties of neural networks (Section 3.2).
3.1 Fundamental Units of Representation
Defining Features.
The notion of a feature in neural networks is a central yet elusive concept, reflecting the pre-paradigmatic state of the field. Traditionally, features are understood as characteristics or attributes of the input data stream (Bishop, 2006). However, a broader interpretation suggests that a feature can be any measurable property or characteristic of a phenomenon, extending beyond human-interpretable elements.
The understanding of features encompasses two perspectives: human-centric and non-human-centric. A human-centric definition posits that features are semantically meaningful, articulable properties of the input, encoded in activation space (Olah, 2022). This view, however, may exclude features that are not understandable to humans. Adversarial examples have been interpreted as evidence for non-interpretable features that are perceptible to neural networks but not to humans (Ilyas et al., 2019). Furthermore, as neural networks evolve to surpass human capabilities, there is no inherent constraint that the features they learn must be comprehensible by humans; instead, they might discover increasingly abstract and alien features (Hubinger, 2019a).
A non-human-centric perspective defines features as independent yet repeatable units that a neural network representation can decompose into (Olah, 2022). This perspective allows for a more comprehensive understanding, encompassing features that are not necessarily interpretable by humans.
@@ 3.1 Fundamental Units of Representation
We adopt the notion of features as the smallest units of how neural networks encode knowledge, such that features cannot be further decomposed into smaller, distinct concepts. These features hypothetically serve as core components of a neural network’s representation, analogous to how cells form the fundamental unit of biological organisms (Olah et al., 2020).
@@
[htb] Definition: Feature Features are the fundamental units of neural network representations.
Figure 2: Comparison of privileged and non-privileged basis in neural networks. Figure adapted from (Bricken et al., 2023).
@@
In the architecture of neural networks, neurons are the natural computational units, potentially representing individual features. Within a neural network representation $h\in\mathbb{R}^{n}$, the $n$ basis directions are called neurons. For a neuron to be meaningful, the basis directions must functionally differ from other directions in the representation, forming a privileged basis – where the basis vectors are architecturally distinguished within the neural network layer from arbitrary directions in activation space. Typical non-linear activation functions privilege the basis directions formed by the neurons, making it meaningful to analyze individual neurons (Elhage et al., 2022b). Analyzing neurons can give insights into a network’s functionality (Sajjad et al., 2022; Mu & Andreas, 2020; Dai et al., 2022; Ghorbani & Zou, 2020; Voita et al., 2023; Durrani et al., 2020; Goh et al., 2021; Bills et al., 2023; Huang et al., 2023).
@@
A neuron corresponding to a single semantic concept is called monosemantic. The intuition behind this term comes from analyzing what inputs activate a given neuron, revealing its associated semantic meaning or concept. If neurons were the fundamental primitives of neural network representations, all neurons would be monosemantic, implying a one-to-one relationship between neurons and features. Comprehensive interpretability would be as tractable as characterizing all neurons and their connections. However, empirically, especially for transformer models (Elhage et al., 2022b), neurons are often observed to be polysemantic, i.e., associated with multiple, unrelated concepts (Arora et al., 2018; Mu & Andreas, 2020; Elhage et al., 2022a; Olah et al., 2020). For example, a single neuron may be activated by both images of cats and images of cars, suggesting it encodes multiple unrelated concepts. Polysemanticity contradicts the interpretation of neurons as fundamental primitives and, in practice, makes it challenging to understand the information processing of neural networks.
@@
  • One trivial scenario would be that feature directions are orthogonal but not aligned with the basis directions (neurons). There is no inherent reason to assume that features would align with neurons in a non-privileged basis, where the basis vectors are not architecturally distinguished. However, even in a privileged basis formed by the neurons, the network could represent features not in the standard basis but as linear combinations of neurons (see Figure 2).
@@
  • Finally, the superposition hypothesis addresses the limitations in the network’s representative capacity — the number of neurons versus the number of crucial concepts. This hypothesis argues that the limited number of neurons compared to the vast array of important concepts necessitates a form of compression. As a result, an $n$-dimensional representation may encode features not with the $n$ basis directions (neurons) but with the $\propto\exp(n)$ possible almost orthogonal directions (Elhage et al., 2022b), leading to polysemanticity.
[htb] Hypothesis: Superposition Neural networks represent more features than they have neurons by encoding features in overlapping combinations of neurons.
@@
Toy models can demonstrate under which conditions superposition occurs (Elhage et al., 2022b; Scherlis et al., 2023). Neural networks, via superposition, may effectively simulate computation with more neurons than they possess by allocating each feature to a linear combination of neurons, creating what is known as an overcomplete linear basis in the representation space. This perspective on superposition suggests that polysemantic models could be seen as compressed versions of hypothetically larger neural networks where each neuron represents a single concept (see Figure 4). Consequently, an alternative definition of features emerges:
[htb] Definition Feature (Alternative) Features are elements that a network would ideally assign to individual neurons if neuron count were not a limiting factor (Bricken et al., 2023). In other words, features correspond to the disentangled concepts that a larger, sparser network with sufficient capacity would learn to represent with individual neurons.
@@
Research on superposition, including works by (Elhage et al., 2022b; Scherlis et al., 2023; Henighan et al., 2023), often investigates simplified models. However, understanding superposition in practical, transformer-based scenarios is crucial for real-world applications, as pioneered by (Gurnee et al., 2023).
@@
This raises the question: How are features encoded if not in discrete neuron units? While a priori features could be encoded in an arbitrarily complex, non-linear structure, a growing body of theoretical arguments and empirical evidence supports the hypothesis that features are commonly represented linearly, i.e., as linear combinations of neurons - hence, as directions in representation space. This perspective promises to enhance our comprehension of neural networks by providing a more interpretable and manipulable framework for their internal representations.
@@
[htb] Hypothesis: Linear Representation Features are directions in activation space, i.e. as linear combinations of neurons.
The linear representation hypothesis is central to neural network analysis, suggesting that networks represent high-level features as linear directions in activation space. This hypothesis simplifies the neural network representations, enhancing their interpretability and ease of manipulation (Nanda et al., 2023b).
The architecture of neural networks, typically comprising linear layers interspersed with non-linear activation functions, inherently favors linear representations. When a neural network layer processes the information from previous layer activations, it typically employs matrix multiplication - only linear features can be processed in a subsequent single linear layer. Conversely, more complex non-linear encodings, though theoretically possible, would require multiple layers to be decoded. Hence, even hypothetical non-linear representations are reducible to intermediate linear representations in the typical neural network design.
An illustrative example can be seen in the work of Li et al. (2023a), who demonstrated that the internal representation of a generative pre-trained transformers (GPT) model trained on the Othello board game could be decoded only with a non-linear probe, not a linear one. This finding suggested that the explicit representation of the board state in terms of "black" and "white" pieces is not present linearly, but other features implicitly represent it. Complementing this, Nanda (2023c); Nanda et al. (2023b) showed that when decoding the board state in terms of "own" and "opponent’s" pieces, a linear probe suffices, thereby reaffirming the linear representation hypothesis. To date, there is no evidence of non-linear representations in neural networks.
@@
Empirical evidence supports the linear representation hypothesis: Firstly, the seminal work by Mikolov et al. (2013) revealing semantic vector calculus in word embeddings pointed to linear representations. Interpretability methods like linear probing (Alain & Bengio, 2016; Belinkov, 2021) and sparse dictionary learning (Bricken et al., 2023; Cunningham et al., 2024; Deng et al., 2023) confirm the linear accessibility of meaningful features. It is possible to decode concepts (O’Mahony et al., 2023), tasks (Hendel et al., 2023), functions (Todd et al., 2023), sentiment (Tigges et al., 2023), and relations (Hernandez et al., 2023; Chanin et al., 2023) linearly in large language models. Additionally, breakthroughs in linear addition for model steering (Turner et al., 2023; Sakarvadia et al., 2023a; Li et al., 2023b) and representation engineering (Zou et al., 2023) highlight the practical implications of linear feature representations regarding model manipulation and interpretability.
@@
While the linear representation hypothesis facilitates interpretability significantly, Sharkey et al. (2022a) warns against neglecting the potential role of non-linear representations. Given the dynamic nature of neural network development, it’s crucial to continuously reevaluate the hypothesis, particularly in light of the possible emergence of non-linear features when interpretability tools that rely on linear representations are subject to optimization pressure (Hubinger, 2022). In this context, the polytope lens provides an alternative perspective, as proposed by Black et al. (2022). This approach shifts the focus to the impact of non-linear activation functions, examining how discrete polytopes, formed by piecewise linear activations, might be the fundamental primitives of neural network representation.
3.2 Computation and Abstraction
@@
[htb] Definition: Circuit Circuits are sub-graphs of the network, consisting of features and the weights connecting them.
@@
The decomposition of neural networks into circuits for interpretability has shown significant promise, particularly in small models trained for specific tasks such as addition, as seen in the work of Nanda et al. (2023a) and Quirke & Barez (2023). However, scaling this analysis to broader behaviors remains challenging. To date, only relatively narrow behaviors like Python docstring formatting (Heimersheim & Jett, 2023) and greater-than-computations (Hanna et al., 2023) have been thoroughly analyzed.
@@
Despite these challenges, there has been notable progress in scaling circuit analysis to larger circuits, such as on GPT-2’s indirect object identification (Wang et al., 2023) and on scaling to larger models such as multiple-choice question answering in Chinchilla (Lieberum et al., 2023). The circuits underlying more general and transferable behaviors are also being explored: McDougall et al. (2023)’s research on copy suppression in GPT-2’s attention heads, for instance, sheds light on model calibration and self-repair mechanisms. Similarly, Davies et al. (2023) and Feng & Steinhardt (2023) focus on how large language models (LLMs) perform variable binding and entity-attribute binding, respectively, providing insights into the representation of symbolic knowledge. Yu et al. (2023) explore mechanisms for factual recall in LLMs, revealing how circuits dynamically balance pre-trained knowledge with new contextual information. Lan & Barez (2023) extend circuit analysis to sequence continuation tasks, identifying shared computational structures across semantically related sequences, thereby enriching our understanding of error identification.
@@
More promisingly, some repeating patterns have shown universality across models and tasks. These universal patterns are called motifs (Olah et al., 2020) and can manifest not just as specific circuits or features but also as higher-level behaviors emerging from the interaction of multiple components. Examples include the curve detectors found across vision models (Cammarata et al., 2021; 2020), induction circuits enabling in-context learning (Olsson et al., 2022), and the phenomenon of branch specialization in neural networks (Voss et al., 2021). Motifs may also capture how models leverage tokens for working memory or parallelize computations in a divide-and-conquer fashion across representations. The significance of motifs lies in revealing the common structures, mechanisms, and strategies that naturally emerge across neural architectures, shedding light on the fundamental building blocks underlying their intelligence.
@@
[htb] Definition: Motif Motifs are repeated patterns within a network, encompassing either features or circuits that emerge across different models and tasks.
@@
Following the evidence for motifs or repeated patterns in neural networks, the universality hypothesis emerges as a pivotal concept. This hypothesis posits a convergence in forming features and circuits across various models and tasks, which could significantly ease interpretability efforts in AI. The universality hypothesis proposes that artificial and biological neural networks share similar features and circuits, suggesting a standard underlying structure (Chan et al., 2023; Sucholutsky et al., 2023; Kornblith et al., 2019). This idea posits that there is a fundamental basis in how neural networks, irrespective of their specific configurations, process and comprehend information. This could be due to inbuilt inductive biases in neural networks or natural abstractions (Chan et al., 2023) – concepts favored by the natural world that any cognitive system would naturally gravitate towards.
[htb] Hypothesis: Universality Analogous features and circuits form across models and tasks (Olah et al., 2020).
Evidence for this hypothesis comes from cross-species neural structures in neuroscience, where similar neural structures and functions are found in different species (Kirchner, 2023). Additionally, machine learning models, including neural networks, tend to converge on similar features, representations, and classifications across different tasks and architectures (Chen et al., 2023a; Hacohen et al., 2020; Li et al., 2015; Bricken et al., 2023). While various studies support the universality hypothesis, questions remain about the extent of feature and circuit similarity across different models and tasks. Nevertheless, this concept bridges AI and other scientific disciplines, offering cross-disciplinary applications and a deeper understanding of artificial and natural cognitive processes. In the context of mechanistic interpretability, this hypothesis has been investigated for neurons (Gurnee et al., 2024), group composition circuits (Chughtai et al., 2023), and modular task processing (Variengien & Winsor, 2023).

### v2

@@ Abstract
Understanding AI systems’ inner workings is critical for ensuring value alignment and safety. This review explores mechanistic interpretability: reverse engineering the computational mechanisms and representations learned by neural networks into human-understandable algorithms and concepts to provide a granular, causal understanding. We establish foundational concepts such as features encoding knowledge within neural activations and hypotheses about their representation and computation. We survey methodologies for causally dissecting model behaviors and assess the relevance of mechanistic interpretability to AI safety. We examine benefits in understanding, control, and alignment, along with risks such as capability gains and dual-use concerns. We investigate challenges surrounding scalability, automation, and comprehensive interpretation. We advocate for clarifying concepts, setting standards, and scaling techniques to handle complex models and behaviors and expand to domains such as vision and reinforcement learning. Mechanistic interpretability could help prevent catastrophic outcomes as AI systems become more powerful and inscrutable.
@@ 1 Introduction
The structure of this paper provides a cohesive overview of mechanistic interpretability, situating the mechanistic approach in the broader interpretability landscape (Section 2), presenting core concepts and hypotheses (Section 3), explaining methods and techniques (Section 4), presenting a taxonomy and survey of the current field (Section 5), exploring relevance to AI safety (Section 6), and addressing challenges (Section 7) and future directions (Section 8).
@@ 2 Interpretability Paradigms from the Outside In
interpretability adopts a top-down approach to unraveling a model’s decision-making processes by probing its learned representations for high-level concepts and patterns governing behavior. Techniques include training supervised auxiliary classifiers (Belinkov, 2021), employing unsupervised contrastive and structured probes (see Section 4.1) to explore latent knowledge (Burns et al., 2023), and using neural representation analysis to quantify the representational similarities between the internal representations learned by different neural networks (Kornblith et al., 2019; Bansal et al., 2021). Beyond observational analysis, concept-based interpretability can enable manipulation of these representations – also called representation engineering (Zou et al., 2023) – potentially enhancing safety by upregulating concepts such as honesty, harmlessness, and morality.
@@ 3 Core Concepts and Assumptions
This section introduces the key concepts and hypotheses of mechanistic interpretability, as summarized in Figure 2. We start by defining features as the basic units of representation (Section 3.1). We then examine the nature of these features, including the challenges posed by polysemantic neurons and the implications of the superposition and linear representation hypotheses (Section 3.2). Next, we explore computation through circuits and motifs, considering the universality hypothesis (Section 3.3). Finally, we discuss the implications for understanding emergent properties, such as internal world models and simulated agents with potentially misaligned objectives (Section 3.4).
Figure 2: Overview of key concepts and hypotheses in mechanistic interpretability, organized into four subsection (pink boxes): defining features (Section 3.1), representation (Section 3.2), computation (Section 3.3), and emergence (Section 3.4). In turquoise, it highlights definitions like features, circuits, and motifs, and in orange, it highlights hypotheses like linear representation, superposition, universality, simulation, and prediction orthogonality. Arrows show relationships, e.g., superposition enabling an alternative feature definition or universality connecting circuits and motifs.
3.1 Defining Features as Representational Primitives
Features as fundamental units of representation.
@@ 3.1 Defining Features as Representational Primitives
The notion of a feature in neural networks is central yet elusive, reflecting the pre-paradigmatic state of mechanistic interpretability. We adopt the notion of features as the fundamental units of neural network representations, such that features cannot be further disentangled into simpler, distinct factors. These features are core components of a neural network’s representation, analogous to how cells form the fundamental unit of biological organisms (Olah et al., 2020).
@@
[htb] Definition 1: Feature Features are the fundamental units of neural network representations that cannot be further decomposed into simpler independent factors.
Concepts as natural abstractions.
The world consists of various entities that can be grouped into categories or concepts based on shared properties. These concepts form high-level summaries like "tree" or "velocity," allowing compact world representations by discarding many irrelevant low-level details. Neural networks can capture and represent such natural abstractions (Chan et al., 2023) through their learned features, which serve as building blocks of their internal representations, aiming to capture the concepts underlying the data.
Features encoding input patterns.
In traditional machine learning, features are understood as characteristics or attributes derived directly from the input data stream (Bishop, 2006). This view is particularly relevant for systems focused on perception, where features map closely to the input data. However, in more advanced systems capable of reasoning with abstractions, features may emerge internally within the model as representational patterns, even when processing information unrelated to the input. In this context, features are better conceptualized as any measurable property or characteristic of a phenomenon (Olah, 2022), encoding abstract concepts rather than strictly reflecting input attributes.
Features as representational atoms.
A key property of features is their irreducibility, meaning they cannot be decomposed into or expressed as a combination of simpler, independent factors. In the context of input-related features, Engels et al. (2024) define a feature as irreducible if it cannot be decomposed into or expressed as a combination of statistically independent patterns or factors in the original input data. Specifically, a feature is reducible if transformations reveal its underlying pattern, which can be separated into independent co-occurring patterns or is a mixture of patterns that never co-occur. We propose generalizing this notion of irreducibility to features encoding abstract concepts not directly tied to input patterns, such that features cannot be reduced to combinations or mixtures of other independent components within the model’s representations.
Features beyond human interpretability.
Features could be defined from a human-centric perspective as semantically meaningful, articulable input patterns encoded in the network’s activation space (Olah, 2022). However, while cognitive systems may converge on similar natural abstractions (Chan et al., 2023), these need not necessarily align with human-interpretable concepts. Adversarial examples have been interpreted as non-interpretable features meaningful to models but not humans. Imperceptible perturbations fool networks, suggesting reliance on alien representational patterns (Ilyas et al., 2019). As models surpass human capabilities, their learned features may become increasingly abstract, encoding information in ways incongruent with human intuition (Hubinger, 2019a). Mechanistic interpretability aims to uncover the actual representations learned, even if diverging from human concepts. While human-interpretable concepts provide guidance, a non-human-centric perspective that defines features as independent model components, whether aligned with human concepts or not, is a more comprehensive and future-proof approach.
3.2 Nature of Features: From Monosemantic Neurons to Non-Linear Representations
Figure 3: Contrasting privileged and non-privileged bases. In a non-privileged basis, there is no reason to expect features to be basis-aligned – calling basis dimensions neurons has no meaning. In a privileged basis, the architecture treats basis directions differently – features can but need not align with neurons (Bricken et al., 2023). Leftmost: Privileged basis; individual features (arrows) align with basis directions, resulting in monosemantic neurons (colored circles). Middle left: Privileged basis, where despite having more features than neurons, some neurons are monosemantic, representing individual features, while others are polysemantic (overlapping gradients), encoding superposition of multiple features. Middle right: Non-privileged basis where, even when the number of features equals the number of neurons, the lack of alignment between the feature directions and basis directions results in polysemantic neurons encoding combinations of features. Rightmost: Non-privileged, polysemantic neurons as feature directions do not align with neuron basis.
@@ 3.2 Nature of Features: From Monosemantic Neurons to Non-Linear Representations
In the architecture of neural networks, neurons are the natural computational units, potentially representing individual features. Within a neural network representation $h\in\mathbb{R}^{n}$, the $n$ basis directions are called neurons. For a neuron to be meaningful, the basis directions must functionally differ from other directions in the representation, forming a privileged basis – where the basis vectors are architecturally distinguished within the neural network layer from arbitrary directions in activation space, as shown in Figure 3. Typical non-linear activation functions privilege the basis directions formed by the neurons, making it meaningful to analyze individual neurons (Elhage et al., 2022b). Analyzing neurons can give insights into a network’s functionality (Sajjad et al., 2022; Mu & Andreas, 2020; Dai et al., 2022; Ghorbani & Zou, 2020; Voita et al., 2023; Durrani et al., 2020; Goh et al., 2021; Bills et al., 2023; Huang et al., 2023).
@@
A neuron corresponding to a single semantic concept is called monosemantic. The intuition behind this term comes from analyzing what inputs activate a given neuron, revealing its associated semantic meaning or concept. If neurons were the representational primitives of neural networks, all neurons would be monosemantic, implying a one-to-one relationship between neurons and features. Comprehensive interpretability would be as tractable as characterizing all neurons and their connections. However, empirically, especially for transformer models (Elhage et al., 2022b), neurons are often observed to be polysemantic, i.e., associated with multiple, unrelated concepts (Arora et al., 2018; Mu & Andreas, 2020; Elhage et al., 2022a; Olah et al., 2020). For example, a single neuron may be activated by both images of cats and images of cars, suggesting it encodes multiple unrelated concepts. Polysemanticity contradicts the interpretation of neurons as representational primitives and, in practice, makes it challenging to understand the information processing of neural networks.
@@
  • One trivial scenario would be that feature directions are orthogonal but not aligned with the basis directions (neurons). There is no inherent reason to assume that features would align with neurons in a non-privileged basis, where the basis vectors are not architecturally distinguished. However, even in a privileged basis formed by the neurons, the network could represent features not in the standard basis but as linear combinations of neurons (see Figure 3, middle right).
@@
  • Finally, the superposition hypothesis addresses the limitations in the network’s representative capacity – the number of neurons versus the number of crucial concepts. This hypothesis argues that the limited number of neurons compared to the vast array of important concepts necessitates a form of compression. As a result, an $n$-dimensional representation may encode features not with the $n$ basis directions (neurons) but with the $\propto\exp(n)$ possible almost orthogonal directions (Elhage et al., 2022b), leading to polysemanticity.
[htb] Hypothesis 1: Superposition Neural networks represent more features than they have neurons by encoding features in overlapping combinations of neurons.
@@
Toy models can demonstrate under which conditions superposition occurs (Elhage et al., 2022b; Scherlis et al., 2023). Neural networks, via superposition, may effectively simulate computation with more neurons than they possess by allocating each feature to a linear combination of neurons, creating what is known as an overcomplete linear basis in the representation space. This perspective on superposition suggests that polysemantic models could be seen as compressed versions of hypothetically larger neural networks where each neuron represents a single concept (see Figure 5). Consequently, an alternative definition of features could be:
[htb] Definition 2: Feature (Alternative) Features are elements that a network would ideally assign to individual neurons if neuron count were not a limiting factor (Bricken et al., 2023). In other words, features correspond to the disentangled concepts that a larger, sparser network with sufficient capacity would learn to represent with individual neurons.
@@
Research on superposition, including works by (Elhage et al., 2022b; Scherlis et al., 2023; Henighan et al., 2023), often investigates simplified models. However, understanding superposition in practical, transformer-based scenarios is crucial for real-world applications, as pioneered by Gurnee et al. (2023).
@@
This raises the question: How are features encoded if not in discrete neuron units? While a priori features could be encoded in an arbitrarily complex, non-linear structure, a growing body of theoretical arguments and empirical evidence supports the hypothesis that features are commonly represented linearly, i.e., as linear combinations of neurons – hence, as directions in representation space. This perspective promises to enhance our comprehension of neural networks by providing a more interpretable and manipulable framework for their internal representations.
@@
[htb] Hypothesis 2: Linear Representation Features are directions in activation space, i.e. linear combinations of neurons.
The linear representation hypothesis suggests that neural networks frequently represent high-level features as linear directions in activation space. This hypothesis can simplify the understanding and manipulation of neural network representations (Nanda et al., 2023b). The prevalence of linear layers in neural network architectures favors linear representations. Matrix multiplication in these layers most readily processes linear features, while more complex non-linear encodings would require multiple layers to decode.
However, recent work by Engels et al. (2024) provides evidence against a strict formulation of the linear representation hypothesis by identifying circular features representing days of the week and months of the year. These multi-dimensional, non-linear representations were shown to be used for solving modular arithmetic problems in days and months. Intervention experiments confirmed that these circular features are the fundamental unit of computation in these tasks, and the authors developed methods to decompose the hidden states, revealing the circular representations.
Establishing non-linearity can be challenging. For example, Li et al. (2023a) initially found that in a GPT model trained on Othello, the board state could only be decoded with a non-linear probe when represented in terms of "black" and "white" pieces, seemingly violating the linearity assumption. However, Nanda (2023c); Nanda et al. (2023b) later showed that a linear probe sufficed when the board state was decoded in terms of "own" and "opponent’s" pieces, reaffirming the linear representation hypothesis in this case. In contrast, the work by Engels et al. (2024) provides a clear and convincing existence proof for non-linear, multi-dimensional representations in language models.
@@
Despite these exceptions, empirical evidence largely supports the linear representation hypothesis in many contexts, especially for feedforward networks with ReLU activations. Semantic vector calculus in word embeddings (Mikolov et al., 2013), successful linear probing (Alain & Bengio, 2016; Belinkov, 2021), sparse dictionary learning (Bricken et al., 2023; Cunningham et al., 2024; Deng et al., 2023), and linear decoding of concepts (O’Mahony et al., 2023), tasks (Hendel et al., 2023), functions (Todd et al., 2023), sentiment (Tigges et al., 2023), and relations (Hernandez et al., 2023; Chanin et al., 2023) in large language models all point to the prevalence of linear representations. Moreover, linear addition techniques for model steering (Turner et al., 2023; Sakarvadia et al., 2023a; Li et al., 2023b) and representation engineering (Zou et al., 2023) highlight the practical implications of linear feature representations.
@@
While the linear representation hypothesis remains a useful simplification, it is important to recognize its limitations and the potential role of non-linear representations (Sharkey et al., 2022a). As neural networks continue to evolve, ongoing reevaluation of the hypothesis is crucial, particularly considering the possible emergence of non-linear features under optimization pressure for interpretability (Hubinger, 2022). Alternative perspectives, such as the polytope lens proposed by Black et al. (2022), emphasize the impact of non-linear activation functions and discrete polytopes formed by piecewise linear activations as potential primitives of neural network representations.
3.3 Circuits as Computational Primitives and Motifs as Universal Circuit Patterns
@@
Figure 6: Comparing observed models (left) and corresponding hypothetical disentangled models (right) trained on similar tasks and data. The observed models show different neuronal activation patterns, while the dissection into feature-level circuits reveals a motif - a shared circuit pattern emerging across models, hinting at universality – models converging on similar solutions based on common underlying principles.
@@
[htb] Definition 3: Circuit Circuits are sub-graphs of the network, consisting of features and the weights connecting them.
@@
The decomposition of neural networks into circuits for interpretability has shown significant promise, particularly in small models trained for specific tasks such as addition, as seen in the work of Nanda et al. (2023a) and Quirke & Barez (2023). Scaling such a comprehensive circuit analysis to broader behaviors in large language models remains challenging. However, there has been notable progress in scaling circuit analysis of narrow behaviors to larger circuits and models, such as indirect object identification (Wang et al., 2023) and greater-than computations (Hanna et al., 2023) in GPT-2 and multiple-choice question answering in Chinchilla (Lieberum et al., 2023).
@@
In search of general and universal circuits, researchers focus particularly on more general and transferable behaviors. McDougall et al. (2023)’s work on copy suppression in GPT-2’s attention heads sheds light on model calibration and self-repair mechanisms. Davies et al. (2023) and Feng & Steinhardt (2023) focus on how large language models represent symbolic knowledge through variable binding and entity-attribute binding, respectively. Yu et al. (2023) explore mechanisms for factual recall, revealing how circuits dynamically balance pre-trained knowledge with new contextual information. Lan & Barez (2023) extend circuit analysis to sequence continuation tasks, identifying shared computational structures across semantically related sequences.
@@
More promisingly, some repeating patterns have shown universality across models and tasks. These universal patterns are called motifs (Olah et al., 2020) and can manifest not just as specific circuits or features but also as higher-level behaviors emerging from the interaction of multiple components. Examples include the curve detectors found across vision models (Cammarata et al., 2021; 2020), induction circuits enabling in-context learning (Olsson et al., 2022), and the phenomenon of branch specialization in neural networks (Voss et al., 2021). Motifs may also capture how models leverage tokens for working memory or parallelize computations in a divide-and-conquer fashion across representations. The significance of motifs lies in revealing the common structures, mechanisms, and strategies that naturally emerge across neural architectures, shedding light on the fundamental building blocks underlying their intelligence. Figure 6 contrasts observed neural network models with hypothetical disentangled models, illustrating how a shared circuit pattern can emerge across different models trained on similar tasks and data, hinting at an underlying universality.
@@
[htb] Definition 4: Motif Motifs are repeated patterns within a network, encompassing either features or circuits that emerge across different models and tasks.
@@
Following the evidence for motifs, we can propose two versions for a universality hypothesis regarding the convergence of features and circuits across neural network models:
[htb] Hypothesis 3: Weak Universality There are underlying principles governing how neural networks learn to solve certain tasks. Models will generally converge on analogous solutions that adhere to the common underlying principles. However, the specific features and circuits that implement these principles can vary across different models based on factors like hyperparameters, random seeds, and architectural choices.
[htb] Hypothesis 4: Strong Universality The same core features and circuits will universally and consistently arise across all neural network models trained on similar tasks and data distributions and using similar techniques, reflecting a set of fundamental computational motifs that neural networks inherently gravitate towards when learning.
The universality hypothesis posits a convergence in forming features and circuits across various models and tasks, which could significantly ease interpretability efforts in AI. It proposes that artificial and biological neural networks share similar features and circuits, suggesting a standard underlying structure (Chan et al., 2023; Sucholutsky et al., 2023; Kornblith et al., 2019). This idea posits that there is a fundamental basis in how neural networks, irrespective of their specific configurations, process and comprehend information. This could be due to inbuilt inductive biases in neural networks or natural abstractions (Chan et al., 2023) – concepts favored by the natural world that any cognitive system would naturally gravitate towards.
Evidence for this hypothesis comes from cross-species neural structures in neuroscience, where similar neural structures and functions are found in different species (Kirchner, 2023). Additionally, machine learning models, including neural networks, tend to converge on similar features, representations, and classifications across different tasks and architectures (Chen et al., 2023a; Hacohen et al., 2020; Li et al., 2015; Bricken et al., 2023). Marchetti et al. (2023) provide mathematical support for emerging universal features.
While various studies support the universality hypothesis, questions remain about the extent of feature and circuit similarity across different models and tasks. In the context of mechanistic interpretability, this hypothesis has been investigated for neurons (Gurnee et al., 2024), group composition circuits (Chughtai et al., 2023), and modular task processing (Variengien & Winsor, 2023), with evidence for the weak but not the strong formulation (Chughtai et al., 2023).
3.4 Emergence of World Models and Simulated Agents